# Computing Richelot isogeny graph of superspecial abelian threefolds

### v1

@@ 3.4. Algorithm for computing the graph $\mathcal{G}_{3}(2,p)$
Apply Thomae’s formula (cf. [32, Theorem 8.1]) for $g=1$. ∎
[…]
@@
[…]
In Steps 15–35, we compute each of the $135$ edges starting from a vertex in turn. If the codomain corresponds to a hyperelliptic curve of genus $3$, then we compute its Shioda invariant (cf. [37]). If the codomain corresponds to a plane quartic, we compute its Dixmier-Ohno invariant (cf. [12] and [33]). These invariants are given by the polynomials whose degrees are bounded by a constant, and hence they can be computed in constant time.
@@
For plane quartics $C,C^{\prime}$ over a field of characteristic 00, it is shown that $C\cong C^{\prime}$ if and only if their Diximier-Ohno invariants are equal to each other. For ones over a field of positive characteristic this is not proven, but it is likely to work also in characteristic $p>7$, see [29] for details.
If $[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i}$ is corresonding to an abelian threefold $A$ of Type $3$, we need to recover an elliptic curve $E$ and a genus-2 curve $C$ such that $A\cong E\times{\rm Jac}(C)$. Let $S_{\rm van}$ be the set of indices $i\in I_{\rm even}$ such that ${\vartheta^{\prime}}_{\!\!i}^{2}=0$, then it follows from Proposition 2.15 that we have that #$\#S_{\rm van}=6$. If $S_{\rm van}=\{27,31,45,47,59,61\}$, then we can restore theta null-points of $E$ and ${\rm Jac}(C)$ by using Lemma 2.8. Even if not, this can be done by acting an appropriate $M\in{\rm Sp}_{6}(\mathbb{Z})$ on $[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i}$ to satisfy $S_{\rm van}=\{27,31,45,47,59,61\}$, and thus we can do computations in Step 28. Also in the case that $A$ is of Type $4$, acting $M\in{\rm Sp}_{6}(\mathbb{Z})$ so that $S_{\rm van}=\{27,31,45,47,54,55,59,61,62\}$ satisfies, we can do computations in Step 31.
Remark 3.12.
In Step 34, we add $(inv,inv^{\prime})$ to the list $\mathcal{E}$ if there exists an edge from the vertex with the invariant $inv$ to the vertex with the invriant $inv^{\prime}$. If we want to count edges with multiplicities, all we need to do is change $\mathcal{E}$ to the list with mulitplicity.
Summing up the above discussions, we obtain Algorithm 3.7. We remark that our algorithm is still valid over a field of characteristic $p\geq 11$. In the following, we assume the following assertions:
Assumption 3.13.
Let $[\vartheta_{i}^{2}]_{i}\in\mathbb{P}^{63}$ be a theta null-points of a superspecial abelian threefold. Then, we can take each $\vartheta_{i}$ such that $\vartheta_{i}\in\mathbb{F}_{p^{2}}$.
We have confirmed that this assertion is correct for all $p<100$. Moreover, we have also experimented with some superspecial abelian threefolds for $p<10000$, and no counterexamples have been found to date.
@@
Under the assumption stated above, we see that all computations in Algorithm 3.7 can be done over $\mathbb{F}_{p^{2}}$. It follows from
@@
that we can do Steps 2–6 within the complexity $\widetilde{O}(p^{3})$, and Steps 7–11 within the complexity $\widetilde{O}(p^{4})$. Next, the loops in Steps 13–37 are done $O(p^{6})$ times. For each loop, Steps 15–35 are executed exactly $135$ times, and moreover all operations can be performed in polynomial time of $\log{p}$. Therefore, Algorithm 3.7 can be done with $\widetilde{O}(p^{6})$ operations in the finite field $\mathbb{F}_{p^{2}}$. ∎
@@ 4.1. Enumeration of superspecial hyperelliptic genus-3 curves
In this subsection, we enumerate superspecial hyperelliptic genus-3 curves in characteristic $p$ with $10011\leq p<$ 100. Indeed, we can do this by modifying Algorithm 3.7 to output the numbers of elements of $\mathcal{V}_{1},\mathcal{V}_{2},\mathcal{V}_{3}$, and $\mathcal{V}_{4}$. Here, remark that #$\#\mathcal{V}_{1}$ (resp. #$\#\mathcal{V}_{2}$) is equal to the number of isomorphism classes of superspecial non-hyperelliptic (resp. hyperelliptic) curves of genus $3$.
@@
As there is no need to output the list $\mathcal{E}$, and therefore Steps $16,18$, and $20$ can be ignored. Moreover, the condition $k\hskip 0.56905pt>\#\mathcal{S}$ in Step $23$ can be replaced by the condition #$\#\mathcal{V}_{1}+\#\mathcal{V}_{2}=\#\Lambda_{3,p}$, where #$\#\Lambda_{3,p}$ is the number of superspecial genus-3 curves (this number is obtained from [3, Theorem 3.10]). However, for the purpose of verifying the validity of our algorithm, we executed it without changing the condition $k\hskip 0.56905pt>\#\mathcal{S}$, in the following experiment.
The following table shows the cardinalities of $\mathcal{V}_{1},\mathcal{V}_{2},\mathcal{V}_{3},\mathcal{V}_{4}$ and the time it took. Here, the column total (resp. compl) denotes the number of times Step 13–37 are performed until the condition $k\hskip 0.56905pt>\#\mathcal{S}$ (resp. #$\#\mathcal{V}_{1}+\#\mathcal{V}_{2}=\#\Lambda_{3,p}$) is satisfied. We note that ${\tt total}$ is equal to the sum of #$\#\mathcal{V}_{1},\#\mathcal{V}_{2},\#\mathcal{V}_{3}$, and #$\#\mathcal{V}_{4}$.
char.   #$\#\mathcal{V}_{1}$   #$\#\mathcal{V}_{2}$   #$\#\mathcal{V}_{3}$   #$\#\mathcal{V}_{4}$ total compl Time (sec)
@@
Table 1. The number of elements of lists $\mathcal{V}_{1},\mathcal{V}_{2},\mathcal{V}_{3}$ and $\mathcal{V}_{4}$
Notably, the final results obtained show that #$\#\mathcal{V}_{1}+\#\mathcal{V}_{2}=\#\Lambda_{3,p}$, and therefore our result is consistent with Brock’s result [3, Theorem 3.10].
@@
For the cardinalities of $\mathcal{V}_{3}$ and $\mathcal{V}_{4}$, we have
#$\#\mathcal{V}_{3}=\#\Lambda_{1,p}\cdot\#\Lambda_{2,p},\quad\#\mathcal{V}_{4}=\binom{\#\Lambda_{1,p}+2}{3}.$
As can be seen in Table 1, the ratio ${\tt compl}/{\tt total}$ is approximately equal to $1/3$ for each $p$. Hence, if the termination condition is changed to #$\#\mathcal{V}_{1}+\#\mathcal{V}_{2}=\#\Lambda_{3,p}$, and then the time required for computing #$\#\mathcal{V}_{2}$ is considered to be only $1/3$ of the time written in Table 1. In that case, listing up all superspecial genus-3 curves in characteristic $p=97$ takes about a day.
4.2. The existence of a superspecial hyperellitptic genus-3 curve
In this subsection, we check that there exists at least one superspecial hyperelliptic curve of genus $3$ in characteristic $p$ with $100007\leq p<$ 10000. First, the sufficient condition for the characteristic $p$ for such a curve to exist is given by the following proposition:
Theorem 4.3 ([3, Theorem 3.15] or [40, Theorem 2]).
@@ 4.2. The existence of a superspecial hyperellitptic genus-3 curve
By the above theorem, there exists a superspecial hyperelliptic curve of genus $3$ if $p\equiv 6\pmod{7}$ or $p\equiv 3\pmod{4}$. To investigate the other cases, we implemented the probabilistic algorithm (Algorithm 4.8) which outputs such a curve modifying Algorithm 3.7 slightly:
@@
4:Compute a theta null-point $[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i}$ of $E^{3}$ by using Lemma 3.9 and Lemma 2.8
@@
We can obtain the defining equation of a superspecial hyperelliptic genus-3 curve in the following procedure:
@@
We remark that Algorithm 4.8 may return different outputs depending on randomness in Steps $1$ and $4$.
@@ 4.3. Drawing the graph $\mathcal{G}_{3}(2,p)$ for small $p$
As a direct result obtained from executing main algorithm, we draw the graph $\mathcal{G}_{3}(2,p)$ for $11\leq p\leq 19$ as follows:
Figure 1. The graph $\mathcal{G}_{3}(2,11)$
@@ 5. Concluding remarks
The complexity of our algorithm is estimated as $\widetilde{O}(p^{6})$ operations in the field $\mathbb{F}_{p^{2}}$ under Remark 3.11 and Assumption 3.13, but these are open problems. Therefore, theoretical proofs of them are future works.
@@
Under this expectation, the complexity of Algorithm 4.8 can be estimated as $\widetilde{O}(p^{2})$ operations in the field $\mathbb{F}_{p^{2}}$. A theoretical proof of the above expectation is also one of our future works.
Ryo Ohashi: Graduate School of Information Science and Technology, The University of Tokyo — 7-3-1 Hongo, Bunkyo-ku, Tokyo, 113-0033, Japan.
Hiroshi Onuki: Graduate School of Information Science and Technology, The University of Tokyo — 7-3-1 Hongo, Bunkyo-ku, Tokyo, 113-0033, Japan.
Momonari Kudo: Fukuoka Institute of Technology — 3-30-1 Wajiro-higashi, Higashi-ku, Fukuoka, 811-0295, Japan.
E-mail address: m-kudo@fit.ac.jp.
Ryo Yoshizumi: Joint Graduate Program of Mathematics for Innovation, Kyushu University — 744 Motooka, Nishi-ku, Fukuoka 819-0395, Japan.
Koji Nuida: Institute of Mathematics for Industry, Kyushu university — 744 Motooka, Nishi-ku, Fukuoka 819-0395, Japan / National Institute of Advanced Industrial Science and Technology (AIST) — 2-3-26 Aomi, Koto-ku, Tokyo, 135-0064, Japan.
E-mail address: nuida@imi.kyushu-u.ac.jp.

### v2

@@ 3.4. Algorithm for listing superspecial genus-3 curves
First, recall from [45, Corollary 2.11.4] that a theta null-point of a principally polarized abelian variety $A$ of dimension $g$ is $\mathbb{F}_{p^{2}}$-rational if and only if there exists a symplectic basis $(e_{1},\ldots,e_{g},{e^{\prime}}_{1},\ldots,{e^{\prime}}_{g})\in A[2]$ such that
  • (i)
    all $e_{i},{e^{\prime}}_{i}$ are $\mathbb{F}_{p^{2}}$-rational, and
  • (ii)
    $e_{T,2}(e_{i},e_{i})=e_{T,2}({e^{\prime}}_{i},{e^{\prime}}_{i})=1$ where $e_{T,2}$ denotes the $2$-Tate pairing.
Here, it is well-known (cf. [29, Theorem 2.6]) that a superspecial curve $C$ descends to a curve $C^{\prime}$ over $\mathbb{F}_{p^{2}}$ where the $p^{2}$-power Frobenius map $\pi_{p^{2}}$ is equal to ±$\pm p$. Then, it follows from Theorem 2.9 that it suffices to show that the theta null-point of ${\rm Jac}(C^{\prime})$ is $\mathbb{F}_{p^{2}}$-rational, since $\sqrt{-1}$ belongs to $\mathbb{F}_{p^{2}}$.
By Ekedahl’s result [14, Proposition 1.2], there is an isomorphism ${\rm Jac}(C^{\prime})\cong E^{g}$ over $\mathbb{F}_{p^{2}}$ as a non-polarized abelian variety where $E$ is a supersingular elliptic curve. Since all the $2$-torsion points of $E$ are $\mathbb{F}_{p^{2}}$-rational, all $e_{i},{e^{\prime}}_{i}$ are $\mathbb{F}_{p^{2}}$-rational for any symplectic basis $(e_{1},\ldots,e_{g},{e^{\prime}}_{1},\ldots,{e^{\prime}}_{g})\in{\rm Jac}(C^{\prime})[2]$. Moreover, if we write $2f_{i}=e_{i}$ with $f_{i}\in{\rm Jac}(C^{\prime})[4]$ for each $i\in\{1,\ldots,g\}$, we have that
$e_{T,2}(e_{i},e_{i})$ $=e_{W,2}(e_{i},\pi_{p^{2}}(f_{i})-f_{i})$
$=e_{W,2}(e_{i},(\pm p-1)f_{i})=e_{W,2}(e_{i},e_{i})^{(\pm p-1)/2}=1$
where $e_{W,2}$ denotes the $2$-Weil pairing. In the same way, we obtain $e_{T,2}({e^{\prime}}_{i},{e^{\prime}}_{i})=1$. Since ${\rm Jac}(C^{\prime})$ satisfies the two previous conditions, and hence its theta null-point is $\mathbb{F}_{p^{2}}$-rational. This completes the proof of the theorem. ∎
Corollary 3.15.
The theta null-point $[\vartheta_{0}^{2}:\vartheta_{1}^{2}:\cdots:\vartheta_{63}^{2}]\in\mathbb{P}^{63}$ of any superspecial principally polarized abelian threefold is $\mathbb{F}_{p^{2}}$-rational.
Proof.
Recall from the beginning of Section 2.3 that every superspecial principally polarized abelian threefold is isomorphic to the Jacobian of a superspecial curve or a product of them. In the former case, the theta null-point is $\mathbb{F}_{p^{2}}$-rational by using Theorem 3.14. In the latter case as well, the assertion follows from Lemma 2.8. ∎
Now, we can list all superspecial genus-3 curves in characteristic $p>7$ as in the following Algorithm 3.16 (as inputs $\Lambda_{1,p}$ and $\Lambda_{2,p}$).
1:A prime integer $p>7$ and the two lists $\Lambda_{1,p}$ and $\Lambda_{2,p}$
2:The list $\Lambda_{3,p}$ of isom. classes of s.sp. genus-3 curves
3:Initialize $\mathcal{L}_{1}\leftarrow\emptyset,\,\mathcal{L}_{2}\leftarrow\emptyset$, and $\mathcal{S}\leftarrow\emptyset$
4:Let $\mathcal{L}_{4}$ be the list of isom. classes of $E_{1}\times E_{2}\times E_{3}$ with $E_{1},E_{2},E_{3}\in\Lambda_{1,p}$
5:Let $\mathcal{L}_{3}$ be the list of isom. classes of $E\times{\rm Jac}(C)$ with $E\in\Lambda_{1,p}$ and $C\in\Lambda_{2,p}$
6:Let $\mathcal{S}$ be the list of theta null-points of threefolds in $\mathcal{L}_{3}\sqcup\mathcal{L}_{4}$, and set $k\leftarrow 1$
7:repeat
8:     Let $[\vartheta^{2}_{i}]_{i}$ be the $k$-th theta null-point of $\mathcal{S}$
9:     for all $[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i}\in\hskip 2.27621pt$ComputeAllRichelotIsogenies($[\vartheta^{2}_{i}]_{i}$do
10:         Compute the number $N_{\rm van}$ of indices $i\in I_{\rm even}$ such that ${\vartheta^{\prime}}_{\!\!i}^{2}=0$
11:         if $N_{\rm van}=0$ then
12:              Let $inv^{\prime}$ be the Dixmier-Ohno invariant of ${\sf RestorePlaneQuartic}([{\vartheta^{\prime}}_{\!\!i}^{2}]_{i})$
13:              if $inv^{\prime}\notin\mathcal{L}_{1}$ then
14:                  Add $inv^{\prime}$ to the list $\mathcal{L}_{1}$ and $[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i}$ to the end of $\mathcal{S}$
15:              end if
16:         else if $N_{\rm van}=1$ then
17:              Let $inv^{\prime}$ be the Shioda invariant of ${\sf RestoreHyperellipticCurve}([{\vartheta^{\prime}}_{\!\!i}^{2}]_{i})$
18:              if $inv^{\prime}\notin\mathcal{L}_{2}$ then
19:                  Add $inv^{\prime}$ to the list $\mathcal{L}_{2}$ and $[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i}$ to the end of $\mathcal{S}$
20:              end if
21:         end if
22:     end for
23:     Set $k\leftarrow k+1$
24:until $k\hskip 0.56905pt>\#\mathcal{S}$
25:return  the list of genus-3 curves whose invariants belong to $\mathcal{L}_{1}\sqcup\mathcal{L}_{2}$
Algorithm 3.16  ListUpSspGenus3Curves
In the following, we provide some notes regarding the above Algorithm 3.16.
  • In line 4, we need compute theta null-points of all superspecial principally polarized abelian threefolds of Types 3 and 4. We explain how to get them: The theta null-point of an elliptic curve $y^{2}=x(x-1)(x-t)$ can be given as $[1:\sqrt{t}:\sqrt{1-t}:0]\in\mathbb{P}^{3}$ using Thomae’s formula (cf. [38, Theorem 8.1]). Also, applying formulae in [19, §$\S 7.5$] or [8, §$\S 5.1$], one can compute the theta null-point of the Jacobian of a genus-2 curve. Then, Lemma 2.8 enables us to compute the theta null-point of their product.
@@
To generate the list $\Lambda_{2,p}$, we can also use theta functions (without going into detail), instead of Algorithm 2.21. If we use such a method, then we get theta null-points of all superspecial abelian surfaces in process, and we do not need to use formulas in [19, §$\S 7.5$] or [8, §$\S 5.1$].
  • In lines 7–20, we compute all the $135$ edges starting from a vertex in turn. If the codomain corresponds to an abelian threefold of Type 1, we compute the Dixmier-Ohno invariants (cf. [13][40]) of the underlying plane quartic. If the codomain corresponds to a Type 2 one, then we compute the Shioda invariants (cf. [48]) of the underlying hyperelliptic curve of genus 3. These invariants are given by polynomials of degree bounded by constants, thus these can be computed in constant time. We note that there is no need to do anything for the codomain of Type 3 or 4 since they have already been obtained in lines 2–3.
@@
If smooth plane quartics $C$ and $C^{\prime}$ over an algebraically closed field have different Dixmier-Ohno invariants, then they are not isomorphic to each other. It is proven that the converse holds when the characteristic is 00, and it is suspected that this similarly holds in the case of characteristic $p>7$, see [33] for details. This is the main reason why we assume that $p>7$ in Algorithm 3.16.
  • In line 22, the condition $k\hskip 0.56905pt>\#\mathcal{S}$ can be replaced by #$\#\mathcal{L}_{1}+\#\mathcal{L}_{2}=\#\Lambda_{3,p}$, where #$\#\Lambda_{3,p}$ denotes the number of superspecial genus-3 curves, which can be obtained from Brock [4, Theorem 3.10]. The algorithm terminates faster with the modified condition; however, in our computational experiment, we ran Algorithm 3.16 with the original condition for reasons explained later.
Finally let us show Theorem 1.1, which asserts the complexity of Algorithm 3.16. (even if the condition in line 22 is replaced, the complexity remains the same).
@@
It follows from Corollary 3.15 that all the theta null-points appearing in Algorithm 3.16 are $\mathbb{F}_{p^{2}}$-rational. In lines 7 and 10, it is necessary to take their square roots to obtain theta constants, which belong to $\mathbb{F}_{p^{4}}$. Then, one can see that all the computations in Algorithm 3.16 can be done over $\mathbb{F}_{p^{4}}$.
It follows from
@@
that the complexity of line 2 and line 3 is given by $\widetilde{O}(p^{3})$ and $\widetilde{O}(p^{4})$, respectively. Next, lines 6–20 are repeated $O(p^{6})$ times since ultimately #$\#\mathcal{S}$ equals the number of isomorphism classes of superspecial principally polarized abelian threefolds. For each loop, lines 8–19 are executed exactly $135$ times, and hence all operations can be performed in polynomial time of $\log{p}$. Therefore, Algorithm 3.16 can be done with $\widetilde{O}(p^{6})$ operations over $\mathbb{F}_{p^{4}}$. ∎
@@ 4.1. Enumeration of superspecial hyperelliptic genus-3 curves
In this subsection, we enumerate superspecial hyperelliptic genus-3 curves in characteristics $p$ with $10011\leq p<$ 100. For this, we only need to modify Algorithm 3.16 to output the cardinalities of lists $\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3}$, and $\mathcal{L}_{4}$. We note that #$\#\mathcal{L}_{1}$ (resp. #$\#\mathcal{L}_{2}$) is equal to the number of isomorphism classes of non-hyperelliptic (resp. hyperelliptic) curves of genus 3.
@@
For the cardinalities of $\mathcal{L}_{3}$ and $\mathcal{L}_{4}$, we have
#$\#\mathcal{L}_{3}=\#\Lambda_{1,p}\cdot\#\Lambda_{2,p},\quad\#\mathcal{L}_{4}=\binom{\#\Lambda_{1,p}+2}{3}.\vspace{0.3mm}$
The following table (Table 1) shows the cardinalities of four lists $\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3},\mathcal{L}_{4}$, and the time it took. As mentioned earlier, the time can be shortened by changing the condition $k\hskip 0.56905pt>\#\mathcal{S}$ in line 22 to #$\#\mathcal{L}_{1}+\#\mathcal{L}_{2}=\#\Lambda_{3,p}$. However, for the purpose of verifying that our result are consistent with Brock’s result [4, Theorem 3.10], we executed it without changing the condition in the following experiment:
char.   #$\#\mathcal{L}_{1}$   #$\#\mathcal{L}_{2}$   #$\#\mathcal{L}_{3}$   #$\#\mathcal{L}_{4}$ total compl Time (sec)
@@
Table 1. The cardinalities of lists $\mathcal{L}_{1},\mathcal{L}_{2},\mathcal{L}_{3}$ and $\mathcal{L}_{4}$
Notably, the final results obtained show that #$\#\mathcal{L}_{1}+\#\mathcal{L}_{2}=\#\Lambda_{3,p}$, and therefore our result is consistent with Brock’s result.
Here, the column total (resp. compl) indicates the number of times lines 6–21 are performed until the condition $k\hskip 0.56905pt>\#\mathcal{S}$ (resp. #$\#\mathcal{L}_{1}+\#\mathcal{L}_{2}=\#\Lambda_{3,p}$) is satisfied. Note that ${\tt total}$ is equal to the sum of #$\#\mathcal{L}_{1},\#\mathcal{L}_{2},\#\mathcal{L}_{3}$, and #$\#\mathcal{L}_{4}$. As can be seen in Table 1, the ratio ${\tt compl}/{\tt total}$ is approximately equal to $1/3$ for each $p$. Hence, if the termination condition is changed to #$\#\mathcal{L}_{1}+\#\mathcal{L}_{2}=\#\Lambda_{3,p}$, and then the time required is considered to be only $1/3$ of the time written in Table 1. For example, listing up all superspecial genus-3 curves in characteristic $p=97$ takes about a day.
4.2. The existence of a superspecial hyperellitptic genus-3 curve
In this subsection, we check that there exists at least one superspecial hyperelliptic curve of genus $3$ in characteristic $p$ with $100007\leq p<$ 10000.
@@ 4.2. The existence of a superspecial hyperellitptic genus-3 curve
Moriya-Kudo’s method [36] is more efficient for finding such curves, but their method can only generate curves whose automorphism groups contain the Klein $4$-group. Even if their approach cannot find such a curve for some $p$, there is a possibility that our approach will find them.
First, the sufficient condition for the characteristic $p$ for such a curve to exist is given by the following proposition:
Theorem 4.3 ([4, Theorem 3.15] or [53, Theorem 2]).
@@
By the above theorem, there exists a superspecial hyperelliptic curve of genus $3$ if $p\equiv 6\pmod{7}$ or $p\equiv 3\pmod{4}$. To investigate the other cases, we implemented the probabilistic algorithm (Algorithm 4.4) which outputs such a curve:
@@
4:Compute a theta null-point $[{\vartheta^{\prime}}_{\!\!i}^{2}]_{i}$ of $E^{3}$ by using Lemma 2.8
@@
Then, one can obtain a superspecial hyperelliptic genus-3 curve as follows:
@@
We remark that Algorithm 4.4 may return different outputs depending on randomness in lines $1$ and $4$.
@@ 4.3. Drawing the graph $\mathcal{G}_{3}(2,p)$ for small $p$
By slightly modifying main algorithm, one can draw the graph $\mathcal{G}_{3}(2,p)$ for small $p$ as follows:
@@ 5. Concluding remarks
The complexity of our algorithm is estimated as $\widetilde{O}(p^{6})$ operations in the field $\mathbb{F}_{p^{4}}$. However, in our computational experiments, all the operations were performed over the field $\mathbb{F}_{p^{2}}$ in practice, and hence it may be possible to assert something stronger from Theorem 1.1.
@@
Recall that the order of magnitude of all superspecial genus-3 curves is $O(p^{6})$, then the proportion of hyperelliptic ones among them can be estimated to be about $1/p^{2}$ under this expectation. A theoretical proof of it is also one of our future works.
Acknowledgements
The authors are grateful to Damien Robert and Everett Howe for their helpful comments on earlier version of this paper. This research was conducted under a contract of “Research and development on new generation cryptography for secure wireless communication services” among “Research and Development for Expansion of Radio Wave Resources (JPJ000254)”, which was supported by the Ministry of Internal Affairs and Communications, Japan. This work was also supported by JSPS Grant-in-Aid for Young Scientists 20K14301 and 23K12949, and WISE program (MEXT) at Kyushu University.